Grade 12th passModern Physics

A boat is rowed across a river at the rate of 4.5km/hr. The river flows at the rate of 6km/hr. The velocity of boat in m/s is

Manjeet
7 Years agoGrade 12th pass

1 Answer

Arun
7 Years ago
the boat and river move at a right angle 
so 
4.5^2 + 6^2 = V^2 = 56.25 
V = 7.5 km/hr 
then 
7.5 km/hr * 1000 m/km * 1 hr/3600 s = 2.1 m/s